# BALROG: Benchmarking Agentic LLM and VLM Reasoning On Games

### v1

@@ Appendix C TextWorld
TextWorld [Côté et al., 2019] is a text-based game environment developed by Microsoft Research that allows for the creation and customization of interactive fiction games. In our experiments, we utilize three specific games from the TextWorld domain: “Treasure Hunter”, “The Cooking Game”, and “Coin Collector”. Each task can be generated with different levels of difficulty by changing number of rooms, enabling obstacles and including distractor rooms. We use the generation rules introduced in  Lu et al. [2024].
@@ C.4 TextWorld Results
In Table 8, we provide results for TextWorld. Standard errors are computed using 20 seeds for each of the 3 tasks. GPT-4o once again leads, obtaining more than twice the average progression of its closest competitor Llama 3.1 70B. The coin collector task was by far the most challenging, with GPT-4o managing to solve it only once out of 20 attempts. Gemini models’ APIs often failed to return completions on textworld, flagging the inputs as ”unsafe”, despite there being absolutely no real safety concerns in the textworld gameplays. This made it impossible to complete a full round of evaluation on the Gemini models, thus we marked them as 0% progression.
@@ C.5 Observations
[…]
@@
b) there is a suitable cooking tool in the room, and then use ‘cook . . . with . . . ’ command.
- When you need to chop/slice/dice ingredients, you need to take the knife and the ingredient in your inventory and then ‘slice/chop/dice … with knife’
@@
- When you have all the ingredients (that got processed or cooked according to the menu), you can ‘prepare meal’ in the kitchen and then ‘eat meal’ to win the game.
- The ingredients should EXACTLY match the color in the recipe, but if the recipe doesn’t specify color, any color would be fine. When you ‘take … with …’, use the EXACT name you see.
@@ D.2 Baba Is AI Results
We provide the Baba Is AI results for LLM and VLM mode in Tables 10 and 10. Standard errors are computed using 5 seeds for each of the 40 Baba Is AI tasks. Surprisingly, the Llama models lead, with Llama 3.2 90B surpassing GPT-4o by a 10% margin in language-only mode. Once again, when vision is added, model performance suffers, with only Gemini-1.5-Pro remaining stable.
@@ Appendix E MiniHack
MiniHack [Samvelyan et al., 2021] is a powerful sandbox framework built on top of the NLE [Küttler et al., 2020] that enables researchers to easily design rich and diverse environments for RL. It provides a flexible platform for creating custom RL tasks ranging from simple grid-world navigation to complex, procedurally generated worlds with intricate game mechanics. The framework allows users to define environments using a human-readable description language or a simple Python interface, giving fine-grained control over environment elements such as terrain, objects, monsters, and traps. MiniHack offers a diverse array of tasks, which can be broadly categorized into three main groups: Navigation Tasks, Skill Acquisition Tasks, and Ported Tasks. To enable interaction with language models, we use NetHack Language Wrapper described in the NetHack Appendix F.
@@
From the MiniHack Skill Acquisition Tasks, we picked Quest (with three different difficulty levels, Easy, Medium, and Hard), which challenges the agent to use objects found in the environment to cross a lava river (these objects can provide levitation or freezing abilities), fight monsters, navigate through rooms or mazes and towards the end of the quest use a wand of death to defeat a powerful monster guarding the goal location.
Figure 8: Example of MiniHack Quest Hard task.
@@
We provide MiniHack results for LLM and VLM mode in Tables 12 and 12, standard errors were computed using 5 seeds for each task. Here, GPT-4o and a Gemini-1.5-Pro equal each other both in language-only and vision-language mode, with both models only managing to complete some of the corridor and corridor battle tasks. None of the other models solved any task.
@@ Appendix F NetHack Learning Environment
The NetHack Learning Environment (NLE) [Küttler et al., 2020] is a scalable, procedurally generated, stochastic, rich, and challenging environment designed to drive long-term research in RL on problems such as exploration, planning, skill acquisition, and language-conditioned RL. Built around the classic and highly complex terminal roguelike game NetHack, NLE provides a complex and dynamic environment where agents must navigate through procedurally generated dungeons, interact with hundreds of entity types, and learn to overcome various challenges.
The goal of the player is to descend through procedurally generated dungeon levels while killing monsters, solving puzzles, and gathering better equipment in order to retrieve the Amulet of Yendor and finally ascend back to the surface to win the game. NetHack is notoriously challenging, even for human players. Mastering the game can take years even with online resources like the NetHack Wiki. Success in NetHack demands long-term strategic planning, as a winning game can involve hundreds of thousands of steps, as well as short-term tactics to fight hordes of monsters. Accurate credit assignment is also crucial to understanding which actions contributed to success or failure. NetHack has already been used extensively as a testbed for RL agents [Wołczyk et al., 2024, Piterbarg et al., 2024, Hambro et al., 2022b]; tabula-rasa RL agents particularly struggle due to sparse reward, complex credit assignment, extremely long-time-horizon, and high stochasticity of the game. The current state-of-the-art agent still remains a hand-coded symbolic policy [Hambro et al., 2022a].
@@ F.1 NetHack Language Wrapper
The NetHack Language Wrapper [Goodger et al., 2023] is a tool designed to interface with the NLE and MiniHack by translating non-language observations into text-based representations. This wrapper, converts various NLE observations such as glyphs, blstats, tty_chars, inv_letters, inv_strs, and tty_cursor into readable text equivalents. For example, it transforms the visual display of the game environment into a textual description, including details about the surroundings, inventory, and player statistics. The wrapper also supports text-based actions, allowing users to interact with the environment using commands like wait, apply, and north, which are then converted into the discrete actions required by the NLE. This functionality enables easier interaction with the NetHack environment, particularly for language models.
@@ F.2 New NetHack Progression System
NetHack features an in-game scoring system that rewards players for actions such as killing monsters, identifying objects, eating food, collecting gold and items, and ultimately ascending in the game. However, we argue that this scoring system does not effectively capture true game progression, as players can win the game with scores ranging from a few hundred thousand to several million points. To address this limitation, we developed a novel, data-informed progression metric using a dataset of human-played NetHack games [Hambro et al., 2022b]. Specifically, we recorded the dungeon levels and experience levels achieved in each game, as well as whether the game resulted in an ascension. Utilizing these statistics, we constructed a data-centric progression system where each data point represents the probability of a human player winning the game after reaching a specific dungeon level or experience level. The resulting progression curves are presented in Figure 10. For practical purposes, we define Dungeon Level 1 (Dlvl:1) and Experience Level 1 as representing 0% progression, corresponding to the game’s starting point, and ascension as 100% progression. The agent’s overall progress is thus determined by the highest progression achieved between the dungeon level and experience level attained.
@@ F.3 NetHack Results
Table 13: Language-Only Performance on NLE
@@
llama-3.1-8B-it 0 $\pm$ 0
gemini-1.5-pro 0.31 $\pm$ 0.31
gpt-4o-mini 0 $\pm$ 0
gemini-1.5-flash 0 $\pm$ 0
llama-3.2-90B-it 0 $\pm$ 0
llama-3.2-11B-it 0 $\pm$ 0
llama-3.2-3B-it 0 $\pm$ 0
llama-3.2-1B-it 0 $\pm$ 0
Table 14: Vision-Language Performance on NLE
@@
gpt-4o-mini 0 $\pm$ 0
gemini-1.5-flash 0 $\pm$ 0
llama-3.2-11B-it 0 $\pm$ 0
gemini-1.5-flash 0 $\pm$ 0
llama-3.2-90B-it 0 $\pm$ 0
llama-3.2-11B-it 0 $\pm$ 0
@@ F.4 Observation
Despite having a language wrapper that describes its observations [Goodger et al., 2023], NetHack is not meant to be played with language only, thus we provided the ASCII map in language mode and the RGB tiles map in vision-language mode. In the LLM context, we only keep information that is important to be kept in the long term, i.e., the game message and language observation. Agent stats and inventory are only needed in the current step, so we do not keep them in the context. This is done also to prevent the context length of NetHack to explode out of control.
@@
[…]

### v2

@@ Appendix C TextWorld
TextWorld (Côté et al., 2019) is a text-based game environment developed by Microsoft Research that allows for the creation and customization of interactive fiction games. In our experiments, we utilize three specific games from the TextWorld domain: “Treasure Hunter”, “The Cooking Game”, and “Coin Collector”. Each task can be generated with different levels of difficulty by changing number of rooms, enabling obstacles and including distractor rooms. We use the generation rules introduced in  Lu et al. (2024).
@@ C.4 TextWorld Results
In Table LABEL:textworld_LLM, we provide results for TextWorld. Standard errors are computed using 20 seeds for each of the 3 tasks. GPT-4o once again leads, obtaining more than twice the average progression of its closest competitor Llama 3.1 70B. The coin collector task was by far the most challenging, with GPT-4o managing to solve it only once out of 20 attempts. Gemini models’ APIs often failed to return completions on textworld, flagging the inputs as ”unsafe”, despite there being absolutely no real safety concerns in the textworld gameplays. This made it impossible to complete a full round of evaluation on the Gemini models, thus we marked them as 0% progression.
@@ C.5 Observations
go $<$dir$>$: move the player north, east, south or west.: move the player north, east, south or west. You can only go to directions indicated with an exit or a door.
@@
b) there is a suitable cooking tool in the room, and then use ’cook . . . with . . . ’ command.
- When you need to chop/slice/dice ingredients, you need to take the knife and the ingredient in your inventory and then ’slice/chop/dice … with knife’
@@
- When you have all the ingredients (that got processed or cooked according to the menu), you can ’prepare meal’ in the kitchen and then ’eat meal’ to win the game.
- The ingredients should EXACTLY match the color in the recipe, but if the recipe doesn’t specify color, any color would be fine. When you ’take … with …’, use the EXACT name you see.
@@ D.2 Baba Is AI Results
We provide the Baba Is AI results for LLM and VLM mode in Tables LABEL:babaisai_LLM and LABEL:babaisai_VLM. Standard errors are computed using 5 seeds for each of the 40 Baba Is AI tasks. Surprisingly, the Llama models lead, with Llama 3.2 90B surpassing GPT-4o by a 10% margin in language-only mode. Once again, when vision is added, model performance suffers, with only Gemini-1.5-Pro remaining stable.
@@ Appendix E MiniHack
MiniHack (Samvelyan et al., 2021) is a powerful sandbox framework built on top of the NLE (Küttler et al., 2020) that enables researchers to easily design rich and diverse environments for RL. It provides a flexible platform for creating custom RL tasks ranging from simple grid-world navigation to complex, procedurally generated worlds with intricate game mechanics. The framework allows users to define environments using a human-readable description language or a simple Python interface, giving fine-grained control over environment elements such as terrain, objects, monsters, and traps. MiniHack offers a diverse array of tasks, which can be broadly categorized into three main groups: Navigation Tasks, Skill Acquisition Tasks, and Ported Tasks. To enable interaction with language models, we use NetHack Language Wrapper described in the NetHack Appendix F.
@@
From the MiniHack Skill Acquisition Tasks, we picked Quest (with two different difficulty levels, Easy and Medium), which challenges the agent to use objects found in the environment to cross a lava river (these objects can provide levitation or freezing abilities), fight monsters, navigate through rooms or mazes and towards the end of the quest use a wand of death to defeat a powerful monster guarding the goal location.
Figure 9: Example of MiniHack Quest Easy task.
@@
We provide MiniHack results for LLM and VLM mode in Tables LABEL:minihack_LLM and LABEL:minihack_VLM, standard errors were computed using 5 seeds for each task. Here, GPT-4o and a Gemini-1.5-Pro equal each other both in language-only and vision-language mode, with both models only managing to complete some of the corridor and corridor battle tasks. None of the other models solved any task.
@@ Appendix F NetHack Learning Environment
The NetHack Learning Environment (NLE) (Küttler et al., 2020) is a scalable, procedurally generated, stochastic, rich, and challenging environment designed to drive long-term research in RL on problems such as exploration, planning, skill acquisition, and language-conditioned RL. Built around the classic and highly complex terminal roguelike game NetHack, NLE provides a complex and dynamic environment where agents must navigate through procedurally generated dungeons, interact with hundreds of entity types, and learn to overcome various challenges.
The goal of the player is to descend through procedurally generated dungeon levels while killing monsters, solving puzzles, and gathering better equipment in order to retrieve the Amulet of Yendor and finally ascend back to the surface to win the game. NetHack is notoriously challenging, even for human players. Mastering the game can take years even with online resources like the NetHack Wiki. Success in NetHack demands long-term strategic planning, as a winning game can involve hundreds of thousands of steps, as well as short-term tactics to fight hordes of monsters. Accurate credit assignment is also crucial to understanding which actions contributed to success or failure. NetHack has already been used extensively as a testbed for RL agents (Wołczyk et al., 2024; Piterbarg et al., 2024; Hambro et al., 2022b); tabula-rasa RL agents particularly struggle due to sparse reward, complex credit assignment, extremely long-time-horizon, and high stochasticity of the game. The current state-of-the-art agent still remains a hand-coded symbolic policy (Hambro et al., 2022a).
@@ F.1 NetHack Language Wrapper
The NetHack Language Wrapper (Goodger et al., 2023) is a tool designed to interface with the NLE and MiniHack by translating non-language observations into text-based representations. This wrapper, converts various NLE observations such as glyphs, blstats, tty_chars, inv_letters, inv_strs, and tty_cursor into readable text equivalents. For example, it transforms the visual display of the game environment into a textual description, including details about the surroundings, inventory, and player statistics. The wrapper also supports text-based actions, allowing users to interact with the environment using commands like wait, apply, and north, which are then converted into the discrete actions required by the NLE. This functionality enables easier interaction with the NetHack environment, particularly for language models.
@@ F.2 New NetHack Progression System
NetHack features an in-game scoring system that rewards players for actions such as killing monsters, identifying objects, eating food, collecting gold and items, and ultimately ascending in the game. However, we argue that this scoring system does not effectively capture true game progression, as players can win the game with scores ranging from a few hundred thousand to several million points. To address this limitation, we developed a novel, data-informed progression metric using a dataset of human-played NetHack games (Hambro et al., 2022b). Specifically, we recorded the dungeon levels and experience levels achieved in each game, as well as whether the game resulted in an ascension. Utilizing these statistics, we constructed a data-centric progression system where each data point represents the probability of a human player winning the game after reaching a specific dungeon level or experience level. The resulting progression curves are presented in Figure 11. For practical purposes, we define Dungeon Level 1 (Dlvl:1) and Experience Level 1 as representing 0% progression, corresponding to the game’s starting point, and ascension as 100% progression. The agent’s overall progress is thus determined by the highest progression achieved between the dungeon level and experience level attained.
@@ F.3 NetHack Results
Table 14: LLM Performance on nle
@@
gpt-4o-mini 0.00 $\pm$ 0.00
gemini-1.5-flash 0.00 $\pm$ 0.00
llama-3.2-1b-it 0.00 $\pm$ 0.00
llama-3.2-3b-it 0.00 $\pm$ 0.00
llama-3.1-8b-it 0.00 $\pm$ 0.00
llama-3.2-11b-it 0.00 $\pm$ 0.00
llama-3.2-90b-it 0.00 $\pm$ 0.00
Table 15: VLM Performance on nle
@@
gpt-4o-mini 0.00 $\pm$ 0.00
gemini-1.5-flash 0.00 $\pm$ 0.00
llama-3.2-11b-it 0.00 $\pm$ 0.00
llama-3.2-90b-it 0.00 $\pm$ 0.00
@@ F.4 Observation
Despite having a language wrapper that describes its observations (Goodger et al., 2023), NetHack is not meant to be played with language only, thus we provided the ASCII map in language mode and the RGB tiles map in vision-language mode. In the LLM context, we only keep information that is important to be kept in the long term, i.e., the game message and language observation. Agent stats and inventory are only needed in the current step, so we do not keep them in the context. This is done also to prevent the context length of NetHack to explode out of control.
@@
travel: travel to a specific location on the map (tip: in the next action, specify < or > for stairs, { for fountain, and _ for altar),